The Stewart Company has $1,354,000 in current assets and $555,140 in current liabilities. Its initial inventory level is $297,880, and it will raise funds as additional notes payable and use them to increase inventory. How much can its short-term debt (notes payable) increase without pushing its current ratio below 2.0? Round your answer to the nearest cent.

$  
Answer #1
The current ratio shouldn't go below 2, it means that maximum reduction in current ratio up to 2 is acceptable, so the desired current ratio is 2
Let us assume X be the amount of short term debt (notes payable) which will be used to increase the inventory.
Desired current ratio = ( Current assets + X ) / ( Current liabilities + X )
2 = ( 1354000 + X ) / ( 555140 + X )
2 * ( 555410 + X ) = 1354000 + X
1110820 + 2X = 1354000 + X
2X - X = 1354000 - 1110820
X = 1354000 - 1110820 243180.00